A person walks 33 m East and then walks 38 m at an angle of 35 degrees North of East. What is the magnitude of the total dis-placement ? Answer in units of m.
1
Expert's answer
2020-10-01T10:10:38-0400

Let's assume that person starts at point A, after walking 33m East he/she changes the direction at point B and then stops walking and point C. We have triangular ABC with AB=33m, BC=38m, "\angle B= 35^\circ" .The total displacement is the side of triangular AC. To find it we can use cosine theorem:

"AC^2 = AB^2 + BC^2 - 2 AB \cdot BC \cdot cos35^\circ"

"AC^2 = 2533- 2054=479"

"AC = \sqrt{479}=21.9 \approx 22" m.

Answer: 22 m.